# The Evolution and Internal Structure of Neptunes and Sub-Neptunes

### v1

@@ 3.1 The Importance of Thermal Conductivity
Figure 2: Total conductivity vs.  radius at $t=5$ Gyr for the warm $M_{\text{p}}$=10 M planet. The blue, green, and red curves represent Cond-1, Cond-2, and Cond-3, respectively. The largest contribution to the conductivity is shown by the line styles with solid, dotted, and dashed for radiation (opacity), vibration, and electron-dominated, respectively.
First we compare the different conductivity cases: Cond-1, Cond-2, and Cond-3. Cond-4 will be compared with Cond-1 separately because they only differ in a small region within the planet. We run the evolution for 10 Gyr and compare the three cases. Figure 2 shows the conductivity as a function of the planetary radius at $t=5$ Gyr. We choose this age to better reflect the typical ages of observed exoplanets, in contrast to the final simulation time of $t=10$ Gyr. The qualitative differences between the conductivities remain similar throughout the evolution.
In the outer low-density parts of the planet, the radiative conductivity dominates. In the inner parts, where material is at least partially ionized, most of the thermal conduction is carried by electrons. In the intermediate part, the temperature is not high enough for the material to be ionized, but the material is too dense for photons to carry the energy. In this part, the vibration of the nuclei/molecules takes over. Although it is the most effective means of non-convective energy transport in this temperature-density regime, it is still relatively inefficient. Note that only Cond-1 includes a model for the vibrational part, but assuming a fully ionized electron conductivity, as we did in Cond-2, makes the vibrational part insignificant. In Appendix E we show how the uncertainty in Cond-1 affects the inferred radius of a $M_{\text{p}}$=10 M planet.
In Figure 2, the difference in the outer part of the planet (from $\sim$ 3 R) is caused by the different temperature-density regime of the radiative-opacity tables. The inner part, which is mostly dominated by the electron conductivity, is similar in shape for Cond-1 and Cond-2 but different by roughly one order of magnitude. In Cond-1, the planet has a region where the vibration of the molecules conducts most of the thermal energy. Yet, this mode is relatively inefficient and leads to a relatively low conductivity.
@@
Next, we analyze the temperature profiles at $t=$ 5 Gyr in Figure 3 to explain the general energy transport and its effect on the planetary radius. In this study, the planetary radius is set to be the radius at the 1 bar pressure-level. We also present the conductivity values throughout the interiors. The plots are intended to qualitatively illustrate which parts of the planet cool efficiently and which parts can contract or remain inflated. In Cond-2 the conductivity remains relatively high throughout the planet. Therefore, energy can be transported out of the deep interior efficiently. The large convective zone carries this energy towards the outer envelope and heating it from beneath. While the interior cools, the cooling of the outer envelope is controlled by the efficiency of the radiative cooling through the atmosphere, resulting in a generally hotter outer envelope. Because the outer envelope consists mostly of H-He, which have a high thermal expansion coefficient, the higher temperature envelope remains puffy throughout the entire evolution. On the other hand, in Cond-1 and Cond-3 the energy is efficiently trapped beneath non-convective regions with low conductivity. Both cases have hot interiors with temperatures above $10^{4.4}$K beyond a radius of $r\sim 2$R, which is a much larger region compared to Cond-2. Only a relatively small amount of energy reaches the outer envelope, such that the outer envelope can cool down efficiently and contract, resulting in smaller radii for Cond-1 and Cond-3 in comparison to Cond-2.
@@
Next, we study the evolution of the planets. Figure 4 shows the radius vs. time for three initial energy states and a planet mass of $M_{\text{p}}=10$ M (see Figure 12 in Appendix D for the 55 M and $15$ M mass planets). We find the same phases in the evolution of a planet with a thermal boundary layer as Scheibe et al. (2021). First, the envelope cools down and contracts, depending on the flux emitted throughout the atmosphere. This phase occurs on a timescale of 0.1 - 1 Gyr. Next, the evolution is governed by the flux from the deep interior to the outer envelope throughout conductive layers. The balance between the energy that is released throughout the atmosphere and the energy that is received from the interior determines the size of the planet (i.e., how inflated it is). Then the contraction is mainly determined by the contraction of the deep interior. However, the high-$Z$ material has a relatively low thermal expansion coefficient such that cooling does not lead to large contraction and the planet evolves relatively slowly. In the last phase, the energy reservoir of the deep interior is depleted, and contraction accelerates again.
@@
Cond-1 and Cond-3 are very similar in the early evolution and deviate more as time progresses. Initially, the contraction is very fast until it slows down after $t\approx 2$ Gyr. The initial contraction is caused by the cooling of the outer envelope and stalls as soon as the internal heat prevents further cooling of the envelope. The planetary radius then depends on the heat flux from the deep interior to the outer envelope. Cond-2 is very different from the other two cases and has a larger radius during most of the evolution. Apart from the outer region of the planet, the conductivity in Cond-2 is higher than in the other cases. The higher conductivity leaks more energy from the deep interior to the outer envelope. This keeps the envelope hotter and therefore more extended compared to the cases with lower conductivities. As a result, in this case the envelope is inflated throughout the entire planetary evolution.
@@
Figure 5: The planetary temperature profile at various times. The three different conductivity cases are shown in the panels from left to right for the warm $M_{\text{p}}=10$ M model. The color indicated the planetary age, while the style of the line indicates whether the zone is convective (solid) or non-convective (dotted).
@@
In order to understand which regions can cool and contract and which ones remain inflated, we show the temperature profiles at various times in Figure 5. In Cond-1, the temperature in the deep interior near the center decreases over time. Due to the slow energy transport into the envelope, the envelope contracts and increases the pressure on the layers beneath. This increasing pressure and the low thermal conductivity through the region with the composition gradient cause a slight increase in temperature around the region of $r=3$ R. The higher conductivity of Cond-2, leads to faster cooling of the deep interior. The profiles of the outer envelope change little after $t=0.5$ Gyr. The heat from the interior keeps the envelope more extended earlier than in the other cases.
@@
In Cond-3, the planet cools even slower than Cond-1 in the deep interior. Since the heat remains trapped in the deep interior, the outer envelope can cool down and contract most efficiently in Cond-3. As soon as the energy of the deep interior is depleted in Cond-1 and Cond-2, the planets could contract more, while Cond-3 still fuels its slightly heated envelope. Hence, we expect that if the simulation time were much longer, Cond-3 would lead to an inflated radius over a longer period of time than Cond-1 and Cond-2.
@@ 3.2 A narrow composition transition
Next, we investigate the effect of the conductivity value for a model with a narrow composition transition between the envelope and the interior metallicity. The structure is similar to a differentiated interior with an envelope of H-He and heavy elements on top of a H-He-free interior. Figure 6 shows the radius evolution for the three conductivities cases for the warm 10 M planet. As the evolution occurs faster than for the wide composition transition models, we show a logarithmic time axis for $t\leq 1$ Gyr to emphasize the behavior at early times. The narrow composition region leads to a small region that is dominated by thermal transport in a non-convective region, and therefore, energy is transported much more efficiently.
@@
The high conductivity of Cond-2 allows for a very efficient cooling. The initial cooling phase, where only the envelope contracts, occurs on a timescale of 0.1 Gyr. Before $t=1$ Gyr, most of the internal energy is depleted, such that the flux into the envelope decreases and the planet can contract again. Interestingly, the cold Cond-2 model (dashed blue) expands significantly due to the energy flux from the deep interior. We find that all the initial energy states for Cond-2 converge to a similar radius within 10 Gyr. After that, the contraction is governed by the radiative cooling of the atmosphere.
To better demonstrate the cooling behavior of the interior, Figure 7 shows warm Cond-1 and Cond-2, and the cold Cond-2 temperature-radius profile for various time. Since Cond-3 is similar to Cond-1 it is not shown.
In the case of Cond-1, the sharp composition change leads to a clear layered thermal profile. Two large convective layers are separated by a non-convective layer that is stabilized by the composition gradient. The outer envelope is again non-convective because of efficient energy transport by radiation. After the initial contraction of the outer envelope, the slow-cooling deep interior acts as an energy reservoir that keeps the envelope inflated for Gyrs.
In Cond-2, the thermal energy transport is sufficiently high to decrease the internal energy reservoir by an order of magnitude. After most of the energy is emitted, the outer envelope can cool down and contract much more than in Cond-1 and Cond-3. As mentioned above, the radius of the cold Cond-2 model even increases in the beginning. The early profile at $t=0.01$ Gyr (black lines in Figure 7) has a more contracted envelope than the slightly more evolved profile at $t=0.5$ Gyr. The central temperature drops by almost $5,000$ K within these two time steps, releasing more energy into the outer parts than the atmosphere can emit. At $t\sim 5$ Gyr the temperature gradient across the composition transition approaches zero. The deep interior becomes non-convective as the temperature gradient in the deep interior is lower than the adiabatic gradient. Now, the evolution is governed by the thermal flux that escapes via the planetary atmosphere.
Figure 6: Planetary radius vs. time for a planet with $M_{\text{p}}=10$ M. The model is similar to the one presented in Figure 4 but has a much narrower composition transition ($\Delta q=0.01$). The solid, dashed, and dotted lines correspond to Cond-1, Cond-2, and Cond-3, respectively. The different colors show the different initial energy states. Note that the x-axis is logarithmic between 0.01 and 1 Gyr and linear between 1 and 10 Gyr.
@@ 3.3 Vibrational conductivity model comparison
Next, we compare the two models, Cond-1 and Cond-4, for the vibrational conductivity. Figure 8 shows the total conductivity and the vibrational component at $t=5$ Gyr for the warm $M_{\text{p}}=10$ M planet. In Figure 14 we show the radius over time for Cond-1 and Cond-4. We find only small variations in all the tested masses and initial energy states, where the largest difference is found to be $1.6\%$ for the hot $M_{\text{p}}=10$ M model. The key difference lies in the smaller conductivity in Cond-4. Note that we use the same electronic conductivity in Cond-1 and Cond-4. A different model for the electronic conductivity can increase the region in the planet where the vibrational conductivity dominates, and therefore increase the deviations between Cond-1 and Cond-4. Note that the scaling law assumed in Cond-4 might not be suitable for the density range in the planet, where the vibrational conductivity dominates. This is because in this region, the densities are lower by about a factor of 10 in comparison to the reference conductivity value of $\rho_{\text{ref}}=3.87$ g cm-3. More information on the vibrational conductivity at various conditions is therefore important.
Figure 8: Conductivity vs. radius at $t=5$ Gyr for the warm $M_{\text{p}}=10$ M planet. The blue and brown lines correspond to Cond-1 and Cond-4, respectively. The total conductivity is shown by the solid line line and thick line shows the contribution from the vibrational conductivity. Outside the shown radius window the conductivity is dominated by the electron and radiation contribution.
@@ 3.4 Implications for exoplanet characterization
Figure 9: Planetary radius at $t=5$ Gyr and relative difference between the models and initial entropies. Top: Planetary radius at $t=5$ Gyr for the models with the wide composition gradient. Different markers indicate the different conductivity cases while the different colors correspond to the different initial entropies. The three planet masses $M_{\text{p}}=5$, 10, and 15 M are shown from left to right. Bottom: Relative difference of the inferred radius $\Delta R/R=(R^{\prime}-R)/R$. Green corresponds to the deviation caused by the different assumed conductivities. We compare the warm Cond-1 with the warm Cond-2 and Cond-3. Blue corresponds to the deviation caused by different initial entropy states. We compare the warm Cond-1 with the cold and hot Cond-1.
We find large differences in the inferred radii depending on the model assumptions. This has important implications for the characterization of exoplanets in the relevant mass-regime. Below, we analyze the results for the models with the wide composition gradient. Figure 9 shows the planetary radii at $t=5$ Gyr for the first three conductivity cases, the cold, warm, and hot initial entropy states and the three masses (see Table 2 for the exact values of the radii). The bottom panel shows the relative difference in the inferred radius between when considering different conductivity models and initial entropies. In particular, we compare warm Cond-1 with the warm Cond-2 and warm Cond-3 models. For the comparison in the inferred primordial entropy, we compare the warm Cond-1 with the cold and hot Cond-1 models. Note that all models have the same composition profile and the same bulk composition, yet the evolution leads to very different radii! Therefore, for a given mass, the radius uncertainty is found to be of the order of $\sim 20\%$ depending on the conductivity, and of the order of $\sim 25\%$, depending on the primordial entropy, when considering the spread between the minimum and maximum values in Figure 9. These differences are much larger than the typical measured uncertainty of the planetary radius, which is of the order of $\sim 5\%$, and could go even lower with upcoming missions (e.g., PLATO). As a result, it is clear that theory plays a key role in the characterization of Neptunes and sub-Neptunes and that in order to take full advantage of the data improvements in theory are required. In particular, it is critical to determine the conductivities of various elements at the relevant pressure-temperature regime and to better constrain the post-formation entropy (and composition gradient) of such planets. We hope that future studies will focus on these topics.

### v2

@@ 3.1 The importance of thermal conductivity
Figure 2: Total conductivity vs. radius at $t=5$ Gyr for the warm $M_{\text{p}}$=10 M planet. The blue, green, and red curves represent Cond-1, Cond-2, and Cond-3, respectively. The largest contribution to the conductivity is shown by the line styles, i.e., solid, dotted, and dashed, for radiation (opacity), vibration, and electron dominated, respectively.
We first compared the different conductivity cases: Cond-1, Cond-2, and Cond-3. We compared Cond-4 with Cond-1 separately because they only differ in a small region within the planet. We ran the evolution for 10 Gyr and compared the three cases. Figure 2 shows the conductivity as a function of the planetary radius at $t=5$ Gyr. We chose this age to better reflect the typical ages of observed exoplanets, in contrast to the final simulation time of $t=10$ Gyr. The qualitative differences between the conductivities remained similar throughout the evolution.
In the outer low-density parts of the planet, the radiative conductivity dominates. In the inner parts, where material is at least partially ionized, most of the thermal conduction is carried by electrons. In the intermediate part, the temperature is not high enough for the material to be ionized, but the material is too dense for photons to carry the energy. In this part, the vibration of nuclei and molecules take over. Although it is the most effective means of non-convective energy transport in this temperature-density regime, it is still inefficient. We note that only Cond-1 includes a model for the vibrational part, but assuming a fully ionized electron conductivity, as we did in Cond-2, makes the vibrational part insignificant. In Appendix E we show how the uncertainty in Cond-1 affects the inferred radius of a $M_{\text{p}}$=10 M planet.
In Figure 2, the difference in the outer part of the planet (from $\sim$ 3 R) is caused by the different temperature-density regime of the radiative-opacity tables. The inner part, which is mostly dominated by the electron conductivity, is similar in shape for Cond-1 and Cond-2 but different by roughly one order of magnitude. In Cond-1, the planet has a region where the vibration of the molecules conducts most of the thermal energy. Yet, this mode is inefficient and leads to low conductivity.
@@
Next, we analyze the temperature profiles at $t=$ 5 Gyr in Figure 3 to explain the general energy transport and its effect on the planetary radius. In this study, the planetary radius is set to be the radius at the 1 bar pressure level. We also present the conductivity values throughout the interiors. The plots are intended to qualitatively illustrate which parts of the planet cool efficiently and which parts can contract or remain inflated. In Cond-2 the conductivity remains high throughout the planet. Therefore, energy can be transported out of the deep interior efficiently. The large convective zone carries this energy toward the outer envelope and heats it from beneath. While the interior cools, the cooling of the outer envelope is controlled by the efficiency of the radiative cooling through the atmosphere, resulting in a generally hotter outer envelope. Because the outer envelope consists mostly of H-He, which have a high thermal expansion coefficient, the higher temperature envelope remains puffy throughout the entire evolution. On the other hand, in Cond-1 and Cond-3 the energy is efficiently trapped beneath non-convective regions with low conductivity. Both cases have hot interiors with temperatures above $10^{4.4}$K beyond a radius of $r\sim 2$R, which is a much larger region compared to Cond-2. Only a small amount of energy reaches the outer envelope, and thus the outer envelope can cool down efficiently and contract, resulting in smaller radii for Cond-1 and Cond-3 in comparison to Cond-2.
@@
Next, we study the evolution of the planets. Figure 4 shows the radius against time for three initial energy states and a planet mass of $M_{\text{p}}=10$ M (see Figure 12 in Appendix D for the 55 M and $15$ M mass planets). We find the same phases in the evolution of a planet with a thermal boundary layer as Scheibe et al. (2021). The envelope first cools down and contracts, depending on the flux emitted throughout the atmosphere. This phase occurs on a timescale of 0.1 - 1 Gyr. Next, the evolution is governed by the flux from the deep interior to the outer envelope throughout conductive layers. The balance between the energy that is released throughout the atmosphere and the energy that is received from the interior determines the size of the planet (i.e., how inflated it is). Then the contraction is mainly determined by the contraction of the deep interior. However, the high-$Z$ material has a low thermal expansion coefficient such that cooling does not lead to a large contraction and the planet evolves slowly. In the last phase, the energy reservoir of the deep interior is depleted, and contraction accelerates again.
@@
We highlight that Cond-1 and Cond-3 are very similar during their early evolution and deviate more as time progresses. Initially, the contraction is very fast until it slows down after $t\approx 2$ Gyr. The initial contraction is caused by the cooling of the outer envelope and stalls as soon as the internal heat prevents further cooling of the envelope. The planetary radius then depends on the heat flux from the deep interior to the outer envelope. Cond-2 is very different from the other two cases and has a larger radius during most of the evolution. Apart from the outer region of the planet, the conductivity in Cond-2 is higher than in the other cases. The higher conductivity leaks more energy from the deep interior to the outer envelope. This keeps the envelope hotter and therefore more extended compared to the cases with lower conductivities. As a result, the envelope is inflated throughout the entire planetary evolution in this case.
@@
Figure 5: Planetary temperature profile at various times. The three different conductivity cases are shown in the panels from left to right for the warm $M_{\text{p}}=10$ M model. The color indicates the planetary age, while the style of the line indicates whether the zone is convective (solid) or non-convective (dotted).
@@
In order to understand which regions can cool and contract and which remain inflated, we show the temperature profiles at various times in Figure 5. In Cond-1, the temperature in the deep interior near the center decreases over time. Due to the slow energy transport into the envelope, the envelope contracts and increases the pressure on the layers beneath. This increasing pressure and the low thermal conductivity through the region with the composition gradient cause a slight increase in temperature around the region of $r=3$ R. The higher conductivity of Cond-2 leads to faster cooling of the deep interior. The profiles of the outer envelope change little after $t=0.5$ Gyr. The heat from the interior keeps the envelope more extended earlier than in the other cases.
@@
In Cond-3, the planet cools even slower than Cond-1 in the deep interior. Since the heat remains trapped in the deep interior, the outer envelope can cool down and contract most efficiently in Cond-3. As soon as the energy of the deep interior is depleted in Cond-1 and Cond-2, the planets can contract more, while Cond-3 still fuels its slightly heated envelope. Hence, we expect that if the simulation time were much longer, Cond-3 would lead to an inflated radius over a longer period of time than Cond-1 and Cond-2.
@@ 3.2 A narrow composition transition
Next, we investigate the effect of the conductivity value for a model with a narrow composition transition between the envelope and the interior metallicity. The structure is similar to a differentiated interior with an envelope of H-He and heavy elements on top of a H-He-free interior. Figure 6 shows the radius evolution for the three conductivity cases for the warm 10 M planet. As the evolution occurs faster than for the wide composition transition models, we show a logarithmic time axis for $t\leq 1$ Gyr to emphasize the behavior at early times. The narrow composition region leads to a small region that is dominated by thermal transport in a non-convective region, and therefore, energy is transported much more efficiently.
@@
The high conductivity of Cond-2 allows for very efficient cooling. The initial cooling phase, where only the envelope contracts, occurs on a timescale of 0.1 Gyr. Before $t=1$ Gyr, most of the internal energy is depleted such that the flux into the envelope decreases and the planet can contract again. Interestingly, the cold Cond-2 model (dashed blue) expands significantly due to the energy flux from the deep interior. We find that all the initial energy states for Cond-2 converge to a similar radius within 10 Gyr. After that, the contraction is governed by the radiative cooling of the atmosphere.
To better demonstrate the cooling behavior of the interior, Figure 7 shows the warm Cond-1 and Cond-2, and the cold Cond-2 temperature-radius profile for various times. Since Cond-3 is similar to Cond-1, it is not shown.
In the case of Cond-1, the sharp composition change leads to a clear layered thermal profile. Two large convective layers are separated by a non-convective layer that is stabilized by the composition gradient. The outer envelope is again non-convective because of efficient energy transport by radiation. After the initial contraction of the outer envelope, the slow-cooling deep interior acts as an energy reservoir that keeps the envelope inflated for several gigayears.
In Cond-2, the thermal energy transport is sufficiently high to decrease the internal energy reservoir by an order of magnitude. After most of the energy is emitted, the outer envelope can cool down and contract much more than in Cond-1 and Cond-3. As mentioned above, the radius of the cold Cond-2 model even increases in the beginning. The early profile at $t=0.01$ Gyr (black lines in Figure 7) has a more contracted envelope than the slightly more evolved profile at $t=0.5$ Gyr. The central temperature drops by almost $5,000$ K within these two time steps, releasing more energy into the outer parts than the atmosphere can emit. At $t\sim 5$ Gyr the temperature gradient across the composition transition approaches zero. The deep interior becomes non-convective, as the temperature gradient in the deep interior is lower than the adiabatic gradient. At this point, the evolution is governed by the thermal flux that escapes via the planetary atmosphere.
Figure 6: Planetary radius vs. time for a planet with $M_{\text{p}}=10$ M. The model is similar to the one presented in Figure 4 but has a much narrower composition transition ($\Delta q=0.001$). The solid, dashed, and dotted lines correspond to Cond-1, Cond-2, and Cond-3, respectively. The different colors show the different initial energy states. We note that the x-axis is logarithmic between 0.01 and 1 Gyr and linear between 1 and 10 Gyr.
@@ 3.3 Vibrational conductivity model comparison
Next, we compare the models Cond-1 and Cond-4, which differ only in the treatment of the vibrational conductivity. Figure 8 shows the total conductivity and the vibrational component at $t=5$ Gyr for the warm $M_{\text{p}}=10$ M planet. In Figure 14 we show the radius over time for Cond-1 and Cond-4. We find only small variations in all the tested masses and initial energy states, where the largest difference is found to be $1.6\%$ for the hot $M_{\text{p}}=10$ M model. The key difference lies in the smaller conductivity in Cond-4. We note that we used the same electronic conductivity in Cond-1 and Cond-4. A different model for the electronic conductivity can increase the region in the planet where the vibrational conductivity dominates and therefore increase the deviations between Cond-1 and Cond-4. We also note that the scaling law assumed in Cond-4 might not be suitable for the density range in the planet, where the vibrational conductivity dominates. This is because in this region, the densities are lower by about a factor of ten in comparison to the reference conductivity value of $\rho_{\text{ref}}=3.87$ g cm-3. More information on the vibrational conductivity at various conditions is therefore important.
Figure 8: Conductivity vs. radius at $t=5$ Gyr for the warm $M_{\text{p}}=10$ M planet. The blue and brown lines correspond to Cond-1 and Cond-4, respectively. The total conductivity is shown by the solid line, and the thick line shows the contribution from the vibrational conductivity. Outside the shown radius window, the conductivity is dominated by the electron and radiation contribution.
@@ 3.4 Implications for exoplanet characterization
Figure 9: Planetary radius at $t=5$ Gyr and relative difference between the models and initial entropies. Top: Planetary radius at $t=5$ Gyr for the models with the wide composition gradient. Different markers indicate the different conductivity cases, while the different colors correspond to the different initial entropies. The three planet masses ($M_{\text{p}}=5$, 10, and 15 M) are shown from left to right. Bottom: Relative difference of the inferred radius $\Delta R/R=(R^{\prime}-R)/R$. Green corresponds to the deviation caused by the different assumed conductivities. We compare the warm Cond-1 with the warm Cond-2 and Cond-3. Blue corresponds to the deviation caused by different initial entropy states. We compare the warm Cond-1 with the cold and hot Cond-1.
We found large differences in the inferred radii depending on the model assumptions. This has important implications for the characterization of exoplanets in the relevant mass regime. Below, we analyze the results for the models with the wide composition gradient. Figure 9 shows the planetary radii at $t=5$ Gyr for the first three conductivity cases; the cold, warm, and hot initial entropy states; and the three masses (see Table 2 for the exact values of the radii). The bottom panel shows the relative difference in the inferred radius when considering different conductivity models and initial entropies. In particular, we compare the warm Cond-1 with the warm Cond-2 and warm Cond-3 models. For the comparison in the inferred primordial entropy, we compare the warm Cond-1 with the cold and hot Cond-1 models. We note that all models have the same composition profile and the same bulk composition, yet the evolution still leads to very different radii. Therefore, for a given mass, the radius uncertainty is found to be of the order of $\sim 20\%$, depending on the conductivity, and of the order of $\sim 25\%$, depending on the primordial entropy, when considering the spread between the minimum and maximum values in Figure 9. These differences are much larger than the typical measured uncertainty of the planetary radius, which is of the order of $\sim 5\%$ and could go even lower with upcoming missions (e.g., PLATO). As a result, it is clear that theory plays a key role in the characterization of Neptunes and sub-Neptunes and that in order to take full advantage of the data, improvements in theory are required. In particular, making such improvements is critical to determining the conductivities of various elements at the relevant pressure-temperature regime and to better constraining the post-formation entropy (and composition gradient) of such planets. We hope that future studies will focus on these topics.